# Towards Urban Planing AI Agent in the Age of Agentic AI

### v3

@@ III The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
Building on these capabilities, we conceptualize an agentic AI urban planner paradigm. Figure 1 demonstrates the overall pipeline of an agentic AI urban planner. 1) Task Planner Agent: Given the heterogeneous objectives and constraints, the agentic AI system starts with a Task Planner agent. The task planner agent has three primary objectives. First, it develops a comprehensive understanding of the overall planning task by integrating the diverse requirements and constraints. Since the requirements are highly heterogeneous, the input to the task planner agent might have different forms such as textual documents, images, spatial data, and numerical indicators. This diverse nature necessitates the capability of interpreting and incorporating multimodal inputs. Multimodal abstraction is essential for capturing the dependencies between objectives, constraints, and contextual factors. Second, it decomposes the intricate planning task by adopting a divide-and-conquer paradigm. Instead of generating the planning alternative as a whole in a single step, the task planner agent partitions the process into a series of smaller and more manageable sub-tasks. Each task represents a specific dimension of the planning task and is associated with its own objectives and local constraints, while remaining consistent with the overarching global goals. Third, it makes well-informed decisions on sub-tasks and sub-constraints assignment. The task planner agent determines how each sub-task should be delegated to specialized agents to ensure that the assigned task matches the expertise, computational capabilities, and available domain-specific tools of specialized agents. For instance, given the previous example of redevelopment zone along the Gulf Coast where agentic AI urban planner reconfigures urban layouts toward improved flood resilience in a divide-and-conquer manner, the task planner agent first integrates multimodal inputs, including flood simulations, infrastructure risk maps, and planner-defined intents, to form a comprehensive understanding of the planning context. Based on the understanding, the task planner agent decomposes the multi-objective planning task into a set of tractable sub-tasks, including flood risk modeling, critical infrastructure exposure analysis, NbS scheme generation, and cost–benefit analysis. 2) Specialized Agents: Once the planning task has been decomposed, the resulting sub-tasks are assigned to Specialized Agents for execution. Each specialized agent is dedicated to a clearly defined scope of work, focusing on one particular dimension of the planning task. Instead of addressing the full complexity of the planning task, they operate within the narrower boundaries of their designated sub-tasks, enabling more accurate sub-outputs. Moreover, the specialized agents are empowered with domain-specific methods and practitioner-developed tools so that they have the access to professional expertise and regulatory standards. For instance, in the previous example, it requires 4 specialized agents:
  • Hydrology Agent applies hydrological and hydraulic models (e.g., HEC-RAS, SWMM) to generate flood inundation maps across different return periods
  • Infrastructure Agent employs GIS overlays and facility databases to evaluate the exposure of hospitals, schools, and power stations
  • NbS Agent leverages green infrastructure siting models to identify potential wetland restoration zones, rain garden networks, and green-blue corridors
  • Economics Agent utilizes cost–benefit analysis models to evaluate trade-offs between flood damage reduction and investment cost for each alternative.
3) Fusion Agent: After the specialized agents have completed their corresponding sub-tasks, their outputs are fed forward to the fusion agent. The fusion agent then aggregates the sub-outputs from the specialized agents to avoid potentially conflicting sub-results and synthesize them into a coherent intermediate plan, ensuring internal consistency across spatial, social, and environmental dimensions. By integrating these results, it produces planning alternatives that balance flood protection, infrastructure security, ecological integrity, and financial feasibility. In this way, the fusion agent serves as an intermediate system-level integrator, transforming fragmented sub-outputs into consistent and feasible planning candidates. This agent is critical because specialized agents, while accurate within their own specialized scope, may produce outputs that are misaligned. For instance, an economically optimal solution may conflict with environmental resilience or social equity requirements. In the flooding resilience case, the fusion agent combines flood inundation maps from the hydrology agent, vulnerability assessments from the infrastructure agent, ecological restoration options from the NbS agent and cost–benefit outcomes from the economics agent. 4) Multimodal Generation Agent: Once a set of coherent planning candidates has been synthesized by the fusion agent, a multimodal generation agent eventually transforms the planning candidates into multimodal outputs that are transparent, interpretable and accessible. Instead of solely reporting textual contexts, the multimodal generation agent can produce complementary formats such as narrative summaries, tabular statistics, spatial maps and model configurations, ensuring that both technical experts and non-specialist audiences can effectively engage with the results. For example, in the flooding resilience case, the multimodal generation agent may generate visual flood risk and land-use maps comparing current and improved scenarios, metrics tables highlighting the exposure of critical infrastructure and textual reports comparing alternative NbS schemes (e.g., wetlands vs. rain gardens). In contrast to conventional automated urban planning approaches that often formulate urban planning as a static, one-shot optimization or generation task, the agentic AI paradigm reconceptualizes urban planning as a multi-step, reasoning-driven process. 1) Agentic AI incorporates reasoning and task decomposition to manage the inherent complexity of planning, while previous approaches rely on static and loosely staged generation that fail to capture multi-granularity dynamics and adaptive trade-offs. 2) Agentic AI empowers the specialized agents with domain-specific methods and practitioner-developed tools, ensuring that outputs align with regulatory standards and professional expertise, where traditional models largely ignore decades of planning theory and expert practices. 3) Agentic AI enhances transparency and interpretability by coordinating sub-tasks and integrating them through a fusion agent, while prior approaches often formulate planning as a black-box optimization that obscures trade-offs and undermines legitimacy 4) Agentic AI produces multimodal outputs such as textual reports, tabular summaries, and spatial maps, demonstrating planning alternatives through comprehensive representations and ensuring interpretability, systematic assessment, and practical utility for both experts and the public, whereas traditional approaches typically only yield abstract reconfigurations such as spatial layouts or land-use maps that are mathematically optimized and fails to express the implications of planning decisions from multiple dimensions or perspectives.
IV What to Consider in Learning Urban Planning AI Agents
@@
IV-A Detecting and Prioritizing Real Human Needs Before Urban Planning
@@
Our perspective is to leverage AI to sense and infer urban planning needs from both explicit signals (e.g., surveys, civic platforms) and implicit behaviors (e.g., movement traces, online discourse, built environment scans). For example, by mining public discourse from social media, 311 complaints, and community forums, we can exploit NLP and LLMs to extract topics (e.g., transit dissatisfaction, service breakdowns) and sentiments that reflect resident pain points, detect urgency, importance, locations in user-generated contents, mapping emotional intensity to geographic locations. As another example, by analyzing street imagery, drone footage, or satellite data, we can detect missing sidewalks, poor lighting, slum expansion, or playground deficits, and classify urban elements relevant to walkability, safety, or accessibility. The third example is: by model GPS traces, call detail records, or smartcard data, AI can identify mismatches between population flows and service coverage, such as last-mile connectivity gaps, transit deserts, or areas with extreme commuting burdens.
@@
Finally, geospatial foundation models, pretrained on large-scale maps, satellite images, and terrain data, can infer land-use, infrastructure density, and vulnerability, provide transferrable knowledge to tackle the issues of under-mapped or data-scarce regions.
IV-B Differentiate GenAI in Strategic Macro Planning and Scenario Micro Planning
@@
Strategic planning aims to capture self-organized planning objectives, such as diversity over mixed land uses, connectivity of urban functions, accessibility to transportation, healthcare, utilities, job opportunities, and simulation to verify the optimal planning objectives over long-term dynamics. On the other hand, scenario planning aims to prioritize conditional generation under fine-grained constraints. GenAI models, such as, hierarchical VAEs, conditional diffusion models, and scenario-based reinforcement fine-tuning can empower guided generation of urban planning.
IV-C Human-Machine Collaborative Planning
Urban CoDesign (i.e., human-machine collaborative planning) enabled by conversational generative AI is a paradigm where urban planners and AI systems co-design planning solutions through natural languages and interactive feedback. CoDesign highlights that AI is a responsive partner that generates, explains, and adapts plans based on iterative human input, instead of generating fixed outputs. This is important because urban planning involves complex trade-offs, contextual sensitivity, and normative goals that pure data-driven models often overlook. Feedback from human planners can ground generative planning in policy, social, economic, cultural dimensions.
@@ IV-C Human-Machine Collaborative Planning
IV-D Leveraging Digital Twins For A Close Simulation, Measurement, Planning Loop

### v4

@@ 3.1.3 Domain-specific tool usage
3.2 The Agentic AI Urban Planner Framework
Building on these capabilities, we propose an agentic AI urban planner framework that embeds AI into the planning workflow to produce clearer, more usable planning outputs. Figure 1 demonstrates the overall pipeline of an agentic AI urban planner.
3.2.1 Task Planner Agent
Given the diversity and potential conflicts among planning objectives and constraints, the agentic AI system starts with a Task Planner agent. The task planner agent has three primary functions. First, it builds a comprehensive view of the planning task by integrating diverse requirements and constraints, providing a transparent basis for scenario generation and evaluation. As the inputs span multiple modalities, such as textual documents, images, spatial data, and numerical indicators, the Task Planner interprets and integrates them into a coherent representation. This multimodal fusion ensures a consistent and traceable understanding of context. It also makes explicit the relationships among objectives, constraints, and contextual factors, improving transparency. Second, it breaks the workflow into smaller, manageable steps from the divide-and-conquer perspective. Instead of generating a complete plan in a single step, it structures the work into sub-tasks that remain aligned with the overall goals. Each sub-task targets a specific planning dimension (e.g., land use, mobility, infrastructure) with its own objectives and constraints, while remaining consistent with the overarching goals. Third, it defines and delegates sub-tasks by determining which Specialized Agents should take them on and which professional tools should be applied. This ensures that each task is matched to the right expertise, computational resources, and established planning practices.
3.2.2 Specialized Agents
Once the planning task has been decomposed, the resulting sub-tasks are assigned to Specialized Agents for implementation. Each Specialized Agent is dedicated to a clearly defined scope of work, focusing on one particular dimension of the planning process. Instead of addressing the full complexity of the planning at once, they operate within the narrower boundaries of their designated sub-tasks, generating more accurate and reliable intermediate results. These results enhance the transparency and traceability of the overall process by providing planners with interpretable outputs at each stage. Moreover, Specialized Agents are empowered with domain-specific methods and practitioner-developed tools, enabling them to align outputs with professional expertise and regulatory standards. Taken together, this division of labor creates a complementary set of agents that lays the foundation for integrating results into coherent planning alternatives.
3.2.3 Fusion Agent
After the Specialized Agents have completed their corresponding sub-tasks, their intermediate results are fed forward to the Fusion Agent. The Fusion Agent then aggregates these diverse outputs, resolve potential conflicts, and synthesize them into coherent planning alternatives. In doing so, it ensures internal consistency across spatial, social, environmental, and economic dimensions that would otherwise remain fragmented. Instead of simple aggregation, the Fusion Agent explicitly highlights trade-offs among competing objectives such as cost versus resilience or growth versus equity. By reconciling differences across sub-task outputs, it generates planning candidates that are technically feasible, socially legitimate, and easier for stakeholders to evaluate and compare. This agent is critical because, while Specialized Agents provide accurate results within their narrow scope, those results might be misaligned when considered together. The Fusion Agent addresses this challenge by acting as a system-level integrator, transforming partial results into consistent and actionable planning scenarios.
3.2.4 Multimodal Generation Agent
After the Fusion Agent synthesizes coherent planning candidates, the Multimodal Generation Agent eventually transforms then into multimodal outputs that are transparent, interpretable and accessible to different audiences. Instead of limiting results to technical reports or abstract layouts, this agent produces complementary formats such as narrative summaries, tabular statistics, and spatial maps. These outputs make it easier for professional planners to evaluate trade-offs, while also enabling non-specialist stakeholders to understand and engage with the alternatives. In contrast to conventional automated urban planning approaches that often formulate urban planning as a single optimization step, the agentic AI framework reconceptualizes urban planning as a multi-stage, reasoning-driven process, transforming the overall task from a black-box approach into an adaptive paradigm that is transparent, interpretable, and grounded in professional practice. By organizing conflicting objectives into a structured problem set, the AI system provides urban planners with a transparent overview that clarifies priorities and guides subsequent steps. Building on this foundation, specialized analyses across distinct dimensions draw on methods and tools established by domain experts to deliver credible and interpretable intermediate outputs. The synthesis of these intermediate results produces coherent alternatives that make trade-offs explicit and reinforce decision legitimacy, Finally, multimodal outputs ensure that results remain transparent to experts and accessible to broader stakeholders. This structured pipeline aligns closely with the routine workflow of urban planners, reinforcing clarity at the outset, credibility in intermediate analyses, and accessibility in final outputs. It enables planners to engage the system through natural-language prompts combined with diverse data inputs, iterating with the agents to test ideas and refine directions. By reasoning over complex contexts, the AI system can produce detailed textual narratives alongside preliminary spatial sketches, accelerating early-stage concept development while enriching it with greater depth. In doing so, agentic AI complements professional practice with outputs that are not only technically robust but also imaginative and accessible across multiple modalities. To illustrate its practical value, we align the framework with the routine planning workflow, using flood resilience as a representative example to highlight how agentic AI augments planner-led processes at each stage.
3.3 Flood Resilience Planning: A Motivating Example
Flood resilience planning represents a classic challenge in urban planning because it concentrates many of the tensions that planners face in practice. Objectives are inherently conflicting: protecting property owners often requires costly defenses, while ensuring public safety may demand relocation or restrictions on development. All of this must be managed within strict budget limits. The tools and data required for decision-making are highly fragmented. Hydrological simulations, GIS-based vulnerability assessments, and economic cost–benefit models are usually run individually, making the integration difficult and time-consuming. In traditional planning processes, key trade-offs are typically resolved through expert deliberation and then explained in technical documents, making it challenging for the public to understand or engage with. These difficulties raise a pressing question: how can planners manage conflicting objectives, fragmented tools, and opaque decisions in a way that remains both rigorous and participatory? Agentic AI offers one possible answer. By reasoning through trade-offs, decomposing complex tasks, integrating expert methods, and presenting results in multimodal accessible forms, it can complement existing practice. To illustrate this potential, we trace how the agentic AI pipeline aligns with the familiar stages of the planning workflow, acting like a junior planner.
3.4 Integrating Agentic AI into the Planning Workflow
Urban planners typically follow a structured workflow that starts with problem scoping, followed by collaborative visioning, data collection and analysis, alternative generation, evaluation of trade-offs, and finally public deliberation and recommendation. Each stage has distinct objectives, methods, and outputs, and together they form the backbone of day-to-day planning practice. Building on the agentic AI framework introduced above, we now align its functions with these standard planning stages, showing how it can complement existing practice and address long-standing challenges such as fragmented tools, opaque trade-offs, and limited participation.
3.4.1 Problem Scoping
Problem scoping is the first step in the planning workflow, where planners define the scope of the challenge, clarify objectives, and identify key constraints. At this stage, they collect input from diverse stakeholders—residents, government agencies, and developers, while also aligning with policy documents, regulations, and budget limits. The goal is to establish a shared understanding of the priorities the planning process must address, from balancing flood protection with housing development to weighing costs against public safety. In practice, however, this step is often complicated by fragmented information, inconsistent stakeholder priorities, and limited transparency, since much of the discussion remains within expert groups. Agentic AI can support planners at this stage by integrating diverse input materials such as flood maps, zoning rules, demographic data, and stakeholder goals into a single, coherent problem representation. Beyond simply compiling inputs, the Task Planner Agent highlights where objectives reinforce one another, where they conflict, and which constraints are likely to dominate later stages, establishing a structured problem set that identifies the objectives and constraints requiring attention in subsequent stages. For human planners, this initial structured problem map provides a transparent baseline that clarifies priorities and reveals the tensions that must be carried forward, providing a clearer starting point for decomposing the task into manageable parts and moving into the collaborative visioning process.
3.4.2 Collaborative Visioning
Collaborative visioning follows problem scoping as the stage where planners and stakeholders work together to define common goals and imagine possible futures based on the shared understanding of the overall challenges and constraints. In this step, residents, officials, and developers discuss priorities for the community’s future, often drafting vision statements that capture aspirations for safety, livability, and economic growth. In practice, collaborative visioning can be skewed by imbalances in participation, leading to situations in which dominant perspectives prevail and others are marginalized. As a result, disagreements are frequently left unresolved, which limits the transparency of outputs for subsequent stages. Agentic AI addresses this challenge by analyzing inputs from meetings, documents, and surveys to identify both points of consensus and areas of disagreement, turning vague tensions into explicit trade-offs that enhance transparency and broaden opportunities for engagement. In addition, the Task Planner Agent decomposes broad and often abstract aspirations into more concrete components. For instance, it can separate goals related to safety, livability, and economic growth into distinct components, each of which can be carried forward to specialized analyses. This decomposition not only clarifies priorities at the visioning stage but also ensures that subsequent evaluations remain transparent and traceable across specific dimensions. To support discussion, the Task Planner Agent can also organizes the decomposed goals into a small set of contrasting vision frames (e.g., safety-led, livability-led, growth-led), each with explicit assumptions and indicative targets, providing concrete reference points for dialogue without relying on abstract statements. For planners, this yields a transparent baseline for visioning—one that highlights common ground, makes conflicts visible, and supplies structured materials (e.g., goal matrices and trade-off tables) to guide dialogue and participation.
3.4.3 Data Intake and Analysis
Following problem scoping and visioning, the planning process proceeds to data collection and analysis, where diverse technical tools are applied to build the evidence base for decision-making. Hydrological simulations estimate flood risks, GIS overlays map vulnerable areas, economic models project costs and benefits, and demographic forecasts anticipate population trends. These analyses provide the technical foundation for developing alternatives. However, in practice they are often fragmented, as tools are applied independently, produce results in disparate formats, and lack mechanisms for integration across domains. This fragmentation not only increases the workload for planners but also limits the transparency of how technical evidence is assembled and interpreted. Agentic AI addresses these challenges through Specialized Agents, each dedicated to a specific dimension of the analysis. A hydrology agent, for instance, can run flood models and standardize the results into formats that connect with infrastructure or economic assessments. Similarly, other agents apply GIS methods, economic models, or demographic analyses, while ensuring that their outputs are interpretable and traceable. Since these agents are designed to operate within defined scopes, they can deliver more accurate and consistent intermediate results, reducing duplication and making integration more straightforward. For planners, this stage provides domain-specific results that are accurate, consistent, and traceable. Each output can be examined on its own, with assumptions and methods clearly documented. In this way, planners and stakeholders can verify the validity of each component, ensuring confidence in the evidence base before alternatives are developed.
3.4.4 Alternative Generation
After the data have been analyzed across different domains, the next stage is to generate planning alternatives. In traditional processes, options are assembled from separate technical studies, which often yields inconsistencies across spatial, social, environmental, and economic dimensions. Conflicts among objectives are not systematically articulated, and results lack a common structure for comparison, making rigorous evaluation difficult. Agentic AI strengthen this stage through the Fusion Agent, which synthesizes the standardized outputs from specialized analyses and synthesizes them into coherent planning candidates. By aligning assumptions and scales, the Fusion Agent creates a consistent basis for comparison across domains. It then evaluates each candidate against the stated constraints and makes trade-offs explicit, ensuring that planners can clearly assess the implications of different alternatives. For planners, the outcome is a not a disconnected set of reports but a structured set of comparable alternatives. Each option is presented with its assumptions, constraints, and quantified impacts, while also clarifying trade-offs among objectives. This combination provides a transparent basis for systematic evaluation and supports discussion not only among technical experts and decision-makers but also with the broader public.
3.5 Evaluation & Deliberation
The final stage of the planning workflow is evaluation and deliberation, where proposed alternatives are assessed and discussed by both experts and the public. Traditionally, the findings are communicated through lengthy reports and technical documents. While these formats provide details, they are difficult for non-specialist audiences to interpret, limiting transparency and weakening opportunities for broad participation. Agentic AI addresses this challenge through the Multimodal Generation Agent. Instead of relying on text-heavy documents, it translates analytical results into a range of outputs, including maps, tables, and narrative summaries. These multimodal representations allow experts to examine precise data while enabling broader stakeholders to engage with intuitive visualizations and plain-language descriptions. For planners, this stage produces a set of planning deliverables that present each alternative through multiple, complementary perspectives. Within these deliverables, trade-offs and assumptions are articulated systematically across alternatives, providing a clear and defensible basis for comparison. The results are presented in varied formats such as narrative reports, tables, and maps, which allow experts to examine detailed evidence and enables policymakers and community members to engage with accessible representations. As a result, planners are able to evaluate alternatives in a process that is more transparent, inclusive, and directly connected to decision-making. This example demonstrates that agentic AI can be seamlessly integrated into the standard planning process of urban planners, aligning with each stage of practice rather than replacing it. Like a junior planner, it contributes constructively throughout the workflow by clarifying objectives at the outset, structuring aspirations into manageable components, providing consistent and credible analyses across domains, assembling alternatives into coherent options, and presenting results in accessible forms that support deliberation. Through these capabilities, the agentic AI system enables planners to navigate complexity with greater clarity, evaluate trade-offs with stronger evidence, and communicate outcomes in ways that are more transparent, participatory, and inclusive.
4 What to Consider in Learning Urban Planning AI Agents
@@
4.1 Detecting and Prioritizing Real Human Needs Before Urban Planning
@@
Our perspective is to leverage AI to sense and infer urban planning needs from both explicit signals (e.g., surveys, civic platforms) and implicit behaviors (e.g., movement traces, online discourse, built environment scans). For example, by mining public discourse from social media, 311 complaints, and community forums, we can exploit NLP and LLMs to extract topics (e.g., transit dissatisfaction, service breakdowns) and sentiments that reflect resident pain points, detect urgency, importance, locations in user-generated contents, mapping emotional intensity to geographic locations. As another example, by analyzing street imagery, drone footage, or satellite data, we can detect missing sidewalks, poor lighting, slum expansion, or playground deficits, and classify urban elements relevant to walkability, safety, or accessibility. The third example is: by modeling GPS traces, call detail records, or smartcard data, AI can identify mismatches between population flows and service coverage, such as last-mile connectivity gaps, transit deserts, or areas with extreme commuting burdens.
@@
Finally, geospatial foundation models, pretrained on large-scale maps, satellite images, and terrain data, can infer land-use, infrastructure density, and vulnerability, providing transferrable knowledge to tackle the issues of under-mapped or data-scarce regions.
4.2 Differentiate GenAI in Strategic Macro Planning and Scenario Micro Planning
@@
Strategic planning aims to capture self-organized planning objectives, such as diversity over mixed land uses, connectivity of urban functions, accessibility to transportation, healthcare, utilities, job opportunities, and simulation to verify the optimal planning objectives over long-term dynamics. On the other hand, scenario planning aims to prioritize conditional generation under fine-grained constraints. GenAI models, such as hierarchical VAEs, conditional diffusion models, and scenario-based reinforcement fine-tuning can empower guided generation of urban planning.
4.3 Human-Machine Collaborative Planning
Urban CoDesign (i.e., human-machine collaborative planning) enabled by conversational generative AI is a paradigm where urban planners and AI systems co-design planning solutions through natural language and interactive feedback. CoDesign highlights that AI is a responsive partner that generates, explains, and adapts plans based on iterative human input, instead of generating fixed outputs. This is important because urban planning involves complex trade-offs, contextual sensitivity, and normative goals that pure data-driven models often overlook. Feedback from human planners can ground generative planning in policy, social, economic, cultural dimensions.
@@ 4.3 Human-Machine Collaborative Planning
4.4 Leveraging Digital Twins For A Close Simulation, Measurement, Planning Loop